# Private Evolution Converges

### v1

@@ 7 Conclusion
We identified key algorithmic and analytical issues in prior theory for Private Evolution (PE) and introduced a new theoretical model with formal convergence guarantees. Our version differs from the practical PE in only one step, preserving its core logic and explaining behaviors observed by practitioners. The analysis combines empirical process theory, Wasserstein convergence, and properties of APIs. We also connect PE to the Private Signed Measure Mechanism, offering new insights into its theoretical foundations. Overall, our work significantly deepens the formal understanding of PE.
@@
This work was supported in part by NSF grants CCF-2338772 and CNS-2148359.
@@ Appendix A Related work
Private evolution. There have been multiple follow up works since Lin et al. [31] introduced the framework of Private Evolution. [31, 44] propose PE for images and text using foundation models as APIs. Other works have extended the framework in various ways, including adapting it to the federated setting [26], the tabular data setting [38], allowing access to many APIs [46], and replacing the API with a simulator [30]. In a similar vein, [32] proposed zeroth-order optimization methods based on genetic algorithms before [31] in order to handle non-differentiable accuracy measures. Despite many empirical works on PE, to the best of our knowledge, its theoretical properties are studied only in [31], under a number of unrealistic assumptions, as discussed in Section 3.
@@
Theory of DP synthetic data. The utility of synthetic data is typically analyzed by quantifying how well the data can be used to answer queries. More concretely, let $\mathcal{Q}$ be a set of queries and $q(S)$ the answer of $q\in\mathcal{Q}$ on sensitive dataset $S$. Most of the theory on synthetic DP data measures the quality of a synthetic dataset $S^{\prime}$ with $\max_{q\in\mathcal{Q}}|q(S)-q(S^{\prime})|$[23, 7, 17]. When $\mathcal{Q}$ contains all the bounded Lipschitz queries, then the accuracy measure coincides with $W_{1}(\mu_{S},\mu_{S^{\prime}})$ (see 2.2); this case was studied in [25, 24, 8]. The cases of sparse Lipschitz queries and smooth queries are studied in [15] and [43], respectively. [21] assumes that $S$ is sampled according to an unknown distribution $\mathcal{D}$ and the accuracy measure is given by $\max_{q\in\mathcal{Q}}|\mathbb{E}_{z\sim\mathcal{D}}[q(z)]-q(S^{\prime})|$. In this work, we focus on the setting where the accuracy measure corresponds to the Wasserstein-1 distance, i.e., $\mathcal{Q}$ contains all bounded Lipschitz queries. [3] gives a practical algorithm along with accuracy guarantees for privately answering many queries, but only works for datasets from a finite data space.
@@
Practice of DP synthetic data. We provide a non-exhaustive selection of practical works on DP synthetic data. For a more comprehensive survey, see [20, 12]. Most practical methods consist of training (either from scratch or finetuning) a non-DP generative model with DP-SGD (that is, the gradients are privatized during training). For example, the following methods have been proposed: DP-GAN [45], G-PATE [33], DP Normalizing Flows [42] and DP Diffusion models [13]. [10] uses the Sinkhorn divergence as a loss function and also uses DP-SGD for training. Several works take inspiration from Optimal Transport (in particular, Wasserstein distances). [34] introduces the Smoothed Sliced Wasserstein Distance, making the loss private (instead of the gradients), and the gradient flow associated with this loss was recently studied in [36]. [35] studied how to generate DP synthetic data under Local DP with entropic optimal transport, and is one of the few practical works which comes with theoretical convergence guarantees.
@@
Appendix B Comparison between Algorithms 12 and Practical PE
B.1 Practical Private Evolution
Let us start introducing Algorithm 3, the practical implementation of the PE algorithm from [31].
Algorithm 3 (Practical) Private Evolution [31]
@@ B.1 Practical Private Evolution
B.2 Comparison between Algorithm 1 and practical PE
@@
B.3 More details on the limitations of the analysis of Algorithm 1 from [31]
The theoretical analysis of PE in [31] does not reflect how PE operates in practice. Their convergence proof relies on first showing that noiseless PE (Algorithm 1 with $\sigma=0$) converges, then arguing that when each data point from the sensitive dataset $S$ repeated $B$ times, with high probability the noisy version behaves exactly the same for large $B$.
@@ B.3 More details on the limitations of the analysis of Algorithm 1 from [31]
To prevent noise from overwhelming the vote signal in the noisy case, [31] assumes that each data point in $S$ is repeated $B$ times. This ensures that, with high probability, the NN histogram entries are at least $B/n$ (dominating the added noise) and that the number of times $\mathrm{nint}(n\mu^{\prime}_{t}[i]/B)B$ that $V_{t}[i]$ is included in $S_{t+1}$ is equal to $n\hat{\mu}_{t}[i]$. In other words, with high probability the noisy NN histogram coincides with the non-DP NN histogram, which clearly violates DP when we remove the unrealistic multiplicity assumption. In fact, the exact assumption they make to argue that $S=_{\eta}S_{T}$ is $B\gg\sqrt{d\log(1/\eta)\log(1/\delta)}/\varepsilon$ (Theorem $2$ in [31]). If we let $B$ be a constant, then their convergence proof is valid only for $\varepsilon=\Omega(\sqrt{d\log(1/\eta)\log(1/\delta)})$ or, equivalently, $\eta=\Omega\big{(}e^{-\varepsilon^{2}/(d\log(1/\delta))}\big{)}$.
B.4 Comparison between Algorithms 2 and 1
@@
Appendix C More Definitions
Definition C.1 (Covering and packing number [41]).
@@ Appendix C More Definitions
Definition C.2 (Gaussian and Laplace Complexity).
Let $\cal F$ be a function class on a metric space $(\Omega,\rho)$. Then, the Gaussian complexity [41] of $\mathcal{F}$ is defined as
@@
The Laplacian complexity [24] of $\mathcal{F}$ is defined as
@@
Appendix D Proof of main result
@@
where the inequality follows from $D_{BL}(\hat{\mu}_{t+1},\tilde{\mu}_{t+1})\geq D_{BL}(\tilde{\mu}_{t+1},\mu^{\prime}_{t+1})$ by definition of the projection in $D_{BL}$ distance. Denote by $\mathbb{E}_{t}[\cdot]$ the expectation when conditioning on the randomness up to iteration $t$ of PE. Lemmata D.1 and D.2 give
@@
with $\gamma=\frac{1}{8\pi[(\sqrt{d}+\log(2))^{2}+\log(2)]}$. Further, using the definition of $\operatorname{Variation\_API}(S_{t})$, it is easy to see that $V_{t}$ has $n_{s}(2\lceil\log_{2}(D/\alpha)\rceil+1)$ elements. Letting $V=n_{s}(2\lceil\log_{2}(D/\alpha)\rceil+1)$, Lemma D.3 implies
@@
is the upper bound on $\mathbb{E}[G_{V}({\cal F}_{BL})]$ from Corollary D.1 and
@@
is the upper bound on $\mathbb{E}[W_{1}(\mu^{\prime}_{t+1},\mu_{S_{t+1}})]$ from Lemma D.5. Hence, after $T$ steps of PE we obtain
@@
$\Gamma_{T}\leq\tilde{O}\bigg{(}\frac{\sigma^{1/\max\{d,2\}}}{\gamma}\bigg{)}=\tilde{O}(dD\sigma^{1/d}).$
This proves the first result. For the second, note that if $\varepsilon,\delta<1$, then $\sigma=\frac{4\sqrt{T\log(1.25/\delta)}}{n\varepsilon}$ is enough noise to privatize $T$ iterations of PE by viewing it as the adaptive composition of $T$ Gaussian mechanisms where each of them adds noise to a NN histogram with $\ell_{2}$ sensitivity $\sqrt{2}/n$ ([14], Corollary 3.3). Since we need $T\geq\log(\gamma/\sigma^{1/\max\{d,2\}})/\gamma$, it suffices to set $T=\frac{\log(n\varepsilon/[4\sqrt{\log(1/\delta)}])}{\max\{d,2\}\gamma}=O(\log(n\varepsilon/\sqrt{\log(1/\delta)}))$ to obtain
@@
D.1 Lemmata used in proof of Theorem 4.1
Lemma D.1.
For any $z_{1},z_{2}\in\Omega$, $\mathbb{E}[\min_{z\in\operatorname{Variation\_API}(z_{1})}\|z-z_{2}\|_{2}]\leq(1-\gamma)\|z_{1}-z_{2}\|_{2}+\alpha$ with $\gamma=\frac{1}{8\pi[(\sqrt{d}+\log(2))^{2}+\log(2)]}$, where the expectation is taken w.r.t $\operatorname{Variation\_API}(z_{1})$.
@@
    where we used the lower bound for the expected maximum of Gaussians from [28] and the upper bound for the expected maximum of chi-squared from [9] (Example 2.7), and the facts that $\sigma_{l}=\frac{\alpha 2^{l-1}}{\sqrt{\pi}[(\sqrt{d}+\log(2))^{2}+\log(2)]}$ and $2^{l^{*}-1}\alpha\leq\|z_{1}-z_{2}\|_{2}\leq 2^{l^{*}}\alpha$. Jensen’s inequality together with $\sqrt{1-t}\leq 1-t/2$ for $t\leq 1$ allow us to conclude.
Lemma D.2.
Let $S,S^{\prime}\in\cup_{m\geq 1}\Omega^{m}$ be two datasets, $V=\operatorname{Variation\_API}(S^{\prime})$ and $\hat{\mu}=\operatorname{NN\_histogram}(S,V,\rho)$ (see Algorithm 4). Suppose that for any $z_{1}\in\Omega,z_{2}\in S$, $\mathbb{E}[\min_{z\in\operatorname{Variation\_API}(z_{1})}\rho(z,z_{2})]\leq(1-\gamma)\rho(z_{1},z_{2})+\alpha$. Then,
@@
equals the mass that $\hat{\mu}$ assigns to $V[j]$, since $\hat{\mu}$ comes from a nearest neighbor histogram according to Algorithm 4 ($\hat{\mu}=\operatorname{NN\_histogram}(S,V,\rho)$). Hence, the marginals of $\hat{\pi}$ are $\mu_{S}$ and $\hat{\mu}$, which proves that $\hat{\pi}\in\Pi(\mu_{S},\hat{\mu})$. Next, note that
@@
[…]
@@
[…]
@@
D.2 Upper bounds
We need upper bounds on Gaussian complexity terms and on the convergence of empirical measures in $W_{1}$. Let us start showing an upper bound on the Gaussian complexity via Dudley chaining. This proof follows closely the proof of upper bounds for Laplacian Complexity given in [24].
Lemma D.4 (Gaussian Complexity bound).

### v2

@@ 6 Empirical Results
Experiments on CIFAR-10
Next, we evaluate our theoretical predictions on an image modeling task, using two subclasses (Dog and Plane) of images from the CIFAR-10 dataset cifar10. We closely follow the experimental setup in PE23, using the ImageNet inception embedding szegedy2016rethinking and Improved Diffusion nichol2021improved trained on ImageNet as $\operatorname{Random\_API}$ and $\operatorname{Variation\_API}$ [^4] We use the checkpoint imagenet64_cond_270M_250K.pt from https://github.com/openai/improved-diffusion. Unlike PE23, we create variations by running the diffusion model with different degrees of variation in each iteration, while PE23 fixed the number of variation degrees per iteration and decreased it across iterations. Figure 3 shows that if we run PE with the hyperparameter setting (including $T$ and $n_{s}$) derived from Theorem 4.1, then the FID improves (decreases) with more samples. This empirically supports the main claim of our paper: PE converges with enough samples, as long as hyperparameters are set correctly and with appropriate APIs. We also evaluate performance of PE with different choices of $T$, similar to Figure 2. Figure 4 confirms that using the theoretically-informed number of evolution steps $T$ is effective even on real data. In particular, the experiments on CIFAR10 data presented in PE23 use $T=20$, which turns out to be suboptimal, while $T=\log(n\varepsilon)$ is a better choice.
Figure 3: We set the privacy parameters to $\varepsilon=5,\delta=10^{-4}$. For each $n\in\{100,200,...,600\}$, we set the hyperparameters according to Theorem 4.1 and run PE. We repeat the experiment 33 times and report the average final FID achieved for each value of $n$, and note that it decreases with larger $n$.
Figure 4: We set the privacy parameters to $\varepsilon=5,\delta=10^{-4}$.We consider a fixed number of samples $n=300$. Then, we set the all hyperparameters according to Theorem 4.1, except the number of steps $T$. For each $T\in\{4,8,12,16,20\}$, we run PE 33 times and report the average final FID achieved.
@@ 7 Conclusion
We identified key algorithmic and analytical issues in prior theory for Private Evolution (PE) and introduced a new theoretical model with formal convergence guarantees. Our version preserves the core logic of practical PE and offers a plausible explanation to behaviors observed by practitioners. The analysis combines empirical process theory, Wasserstein convergence, and properties of APIs. We also connect PE to the Private Signed Measure Mechanism, offering new insights into its theoretical foundations. Overall, our work significantly deepens the formal understanding of PE.
Limitations and future work. The choice of $\operatorname{Variation\_API}$ that we presented in Section 3 is not the same as what practical variants of PE utilize. Furthermore, the convergence rate we proved in Theorem 4.1 suffers from the curse of dimensionality, requiring many samples to ensure convergence. Future work could expand the set of theoretically tractable $\operatorname{Variation\_API}$ for which we understand the convergence of PE or study alternative accuracy metrics that do not suffer from the curse of dimensionality.
@@
This work was supported in part by NSF grants CCF-2338772 and CNS-2148359 and the Sloan Foundation. TG was partially funded to attend this conference by the CMU GSA/Provost Conference Funding.
@@ Appendix A Related work
Private evolution. There have been multiple follow up works since Lin et al. [PE23] introduced the framework of Private Evolution. [PE23, pe2] propose PE for images and text using foundation models as APIs. Other works have extended the framework in various ways, including adapting it to the federated setting [hou24PE_federated], the tabular data setting [swanberg2025PEtabulardata], allowing access to many APIs [zou2025PEmultipleLLMs], and replacing the API with a simulator [pe3]. In a similar vein, [liu2023genetic_DPsyntheticdata] proposed zeroth-order optimization methods based on genetic algorithms before [PE23] in order to handle non-differentiable accuracy measures. Despite many empirical works on PE, to the best of our knowledge, its theoretical properties are studied only in [PE23], under a number of unrealistic assumptions, as discussed in Section 3.
@@
Theory of DP synthetic data. The utility of synthetic data is typically analyzed by quantifying how well the data can be used to answer queries. More concretely, let $\mathcal{Q}$ be a set of queries and $q(S)$ the answer of $q\in\mathcal{Q}$ on sensitive dataset $S$. Most of the theory on synthetic DP data measures the quality of a synthetic dataset $S^{\prime}$ with $\max_{q\in\mathcal{Q}}|q(S)-q(S^{\prime})|$[hardt2012simple, blum2013learning, dwork2015efficient]. When $\mathcal{Q}$ contains all the bounded Lipschitz queries, then the accuracy measure coincides with $W_{1}(\mu_{S},\mu_{S^{\prime}})$ (see 2.2); this case was studied in [he2024onlineDPsyntheticdata, Vershynin23DPsyntheticdata, boedihardjo2024metricDP]. The cases of sparse Lipschitz queries and smooth queries are studied in [donhauser2024certified] and [wang2016smooth_queries], respectively. [gonzalez2024mirror] assumes that $S$ is sampled according to an unknown distribution $\mathcal{D}$ and the accuracy measure is given by $\max_{q\in\mathcal{Q}}|\mathbb{E}_{z\sim\mathcal{D}}[q(z)]-q(S^{\prime})|$. In this work, we focus on the setting where the accuracy measure corresponds to the Wasserstein-1 distance, i.e., $\mathcal{Q}$ contains all bounded Lipschitz queries. [aydore21DPQRadaptive_projection] gives a practical algorithm along with accuracy guarantees for privately answering many queries, but only works for datasets from a finite data space.
@@
Practice of DP synthetic data. We provide a non-exhaustive selection of practical works on DP synthetic data. For a more comprehensive survey, see [gong2025dpimagebench, chen2025benchmarking]. Most practical methods consist of training (either from scratch or finetuning) a non-DP generative model with DP-SGD (that is, the gradients are privatized during training). For example, the following methods have been proposed: DP-GAN [xie2018DPgan], G-PATE [long2021gpate], DP Normalizing Flows [DP_normalizingflows] and DP Diffusion models [dockhorn2022dpdifussionmodels]. [cao2021DPGMwithsinkhorn] uses the Sinkhorn divergence as a loss function and also uses DP-SGD for training. Several works take inspiration from Optimal Transport (in particular, Wasserstein distances). [DPslicedWasserstein] introduces the Smoothed Sliced Wasserstein Distance, making the loss private (instead of the gradients), and the gradient flow associated with this loss was recently studied in [sebag2023DPslicedWassersteinflow]. [reshetova2024training_gen_models_with_EOT] studied how to generate DP synthetic data under Local DP with entropic optimal transport, and is one of the few practical works which comes with theoretical convergence guarantees.
@@
Appendix B More simulations
First, we provide the figure that shows how our theory captures the dependence on initialization, as explained in Section 6.
Figure 5: Recall from Section 6 that $\varepsilon=1,\delta=10^{-4}$. We set the number of private samples to $n=1000$. We run PE on different initial datasets created with $\operatorname{Random\_API}$ and plot the accuracy trajectories over $2\log(n\varepsilon)=12$ iterations. We note that the best number of steps depends heavily on the initialization from $\operatorname{Random\_API}$. The private dataset $S$ is a random dataset in $\Omega\cap\mathbb{R}^{2}_{+}$ . Left: When $S_{0}=S$ (i.e., $\Gamma_{0}=0$, the private data is the same as the initial synthetic data), the optimal number of PE steps is 0. Middle: When PE is initialized poorly (e.g., $S_{0}$ consists of only $(0,0)$, so $\Gamma_{0}$ is large), more iterations are needed. Right: Interpolating between the previous cases parametrized by $\beta$: $S_{0}=(1-2\beta)S$, PE can improve or degrade performance depending on $\operatorname{Random\_API}$; it never exceeds the worst-case error bound. Results are averaged over 100 runs. Our analysis explains this phenomenon (see text in Section 6).
B.1 Simulations varying dimension and the privacy parameter
In the simulations presented in Section 6, we fixed the privacy parameters to $\varepsilon=1,\delta=10^{-4}$ and the dimension to $d=2$. We provide simulations that verify two expected behaviors that can be derived from the convergence rate provided in Theorem 4.1 when varying $\varepsilon$ and $d$. Namely, the accuracy degrades as the dimension $d$ grows, and improves when $\varepsilon$ grows.
Figure 6: We set $\delta=10^{-}4$ and run PE with the hyperparameter setting suggested by Theorem 4.1. For the plot on the right, we fixed $\varepsilon=1$ and vary $d\in\{1,2,3,...,10\}$. For each value of $d$, we report the final accuracy of PE averaged over $100$ runs. On the right, we repeat a similar procedure but fixing $d=2$ and varying $\varepsilon\in\{0.1,0.2,...,1\}$.
Appendix C Comparison between Algorithms 12 and Practical PE
C.1 Practical Private Evolution
Let us start introducing Algorithm 3, the practical implementation of the PE algorithm from [PE23].
Algorithm 3 (Practical) Private Evolution [PE23]
@@ C.1 Practical Private Evolution
C.2 Comparison between Algorithm 1 and practical PE
@@
C.3 More details on the limitations of the analysis of Algorithm 1 from [PE23]
The theoretical analysis of PE in [PE23] does not reflect how PE operates in practice. Their convergence proof relies on first showing that noiseless PE (Algorithm 1 with $\sigma=0$) converges, then arguing that when each data point from the sensitive dataset $S$ repeated $B$ times, with high probability the noisy version behaves exactly the same for large $B$.
@@ C.3 More details on the limitations of the analysis of Algorithm 1 from [PE23]
To prevent noise from overwhelming the vote signal in the noisy case, [PE23] assumes that each data point in $S$ is repeated $B$ times. This ensures that, with high probability, the NN histogram entries are at least $B/n$ (dominating the added noise) and that the number of times $\mathrm{nint}(n\mu^{\prime}_{t}[i]/B)B$ that $V_{t}[i]$ is included in $S_{t+1}$ is equal to $n\hat{\mu}_{t}[i]$. In other words, with high probability the noisy NN histogram coincides with the non-DP NN histogram, which clearly violates DP when we remove the unrealistic multiplicity assumption. In fact, the exact assumption they make to argue that $S=_{\eta}S_{T}$ is $B\gg\sqrt{d\log(1/\eta)\log(1/\delta)}/\varepsilon$ (Theorem $2$ in [PE23]). If we let $B$ be a constant, then their convergence proof is valid only for $\varepsilon=\Omega(\sqrt{d\log(1/\eta)\log(1/\delta)})$ or, equivalently, $\eta=\Omega\big(e^{-\varepsilon^{2}/(d\log(1/\delta))}\big)$.
C.4 Comparison between Algorithms 2 and 1
@@
Appendix D More Definitions
Definition D.1 (Covering and packing number [wainwright2019HDS]).
@@ Appendix D More Definitions
Definition D.2 (Gaussian and Laplace Complexity).
[…]
@@
The Laplacian complexity [Vershynin23DPsyntheticdata] of $\mathcal{F}$ is defined as
@@
Appendix E Proof of main result
@@
where the inequality follows from $D_{BL}(\hat{\mu}_{t+1},\tilde{\mu}_{t+1})\geq D_{BL}(\tilde{\mu}_{t+1},\mu^{\prime}_{t+1})$ by definition of the projection in $D_{BL}$ distance. Denote by $\mathbb{E}_{t}[\cdot]$ the expectation when conditioning on the randomness up to iteration $t$ of PE. Lemmata E.1 and E.2 give
@@
with $\gamma=\frac{1}{8\pi[(\sqrt{d}+\log(2))^{2}+\log(2)]}$. Further, using the definition of $\operatorname{Variation\_API}(S_{t})$, it is easy to see that $V_{t}$ has $n_{s}(2\lceil\log_{2}(D/\alpha)\rceil+1)$ elements. Letting $V=n_{s}(2\lceil\log_{2}(D/\alpha)\rceil+1)$, Lemma E.3 implies
@@
is the upper bound on $\mathbb{E}[G_{V}({\cal F}_{BL})]$ from Corollary E.1 and
@@
is the upper bound on $\mathbb{E}[W_{1}(\mu^{\prime}_{t+1},\mu_{S_{t+1}})]$ from Lemma E.5. Hence, after $T$ steps of PE we obtain
@@
$\Gamma_{T}\leq\tilde{O}\bigg(\frac{D\sigma^{1/\max\{d,2\}}}{\gamma}\bigg)=\tilde{O}(dD\sigma^{1/d}).$
This proves the first result. For the second, note that if $\varepsilon,\delta<1$, then $\sigma=\frac{4\sqrt{T\log(1.25/\delta)}}{n\varepsilon}$ is enough noise to privatize $T$ iterations of PE by viewing it as the adaptive composition of $T$ Gaussian mechanisms where each of them adds noise to a NN histogram with $\ell_{2}$ sensitivity $\sqrt{2}/n$ ([dong2022gaussianDP], Corollary 3.3). Since we need $T\geq\log(\gamma/\sigma^{1/\max\{d,2\}})/\gamma$, it suffices to set $T=\frac{\log(n\varepsilon/[4\sqrt{\log(1/\delta)}])}{\max\{d,2\}\gamma}=O(\log(n\varepsilon/\sqrt{\log(1/\delta)}))$ to obtain
@@
E.1 Lemmata used in proof of Theorem 4.1
Lemma E.1.
Assume $\Omega$ is closed and convex. Then, for any $z_{1},z_{2}\in\Omega$, $\mathbb{E}[\min_{z\in\operatorname{Variation\_API}(z_{1})}\|z-z_{2}\|_{2}]\leq(1-\gamma)\|z_{1}-z_{2}\|_{2}+\alpha$ with $\gamma=\frac{1}{8\pi[(\sqrt{d}+\log(2))^{2}+\log(2)]}$, where the expectation is taken w.r.t $\operatorname{Variation\_API}(z_{1})$.
@@
    where we used the lower bound for the expected maximum of Gaussians from [kamath2015boundsmaxgaussians] and the upper bound for the expected maximum of chi-squared from [boucheron2013concentration] (Example 2.7), and the facts that $\sigma_{l}=\frac{\alpha 2^{l-1}}{\sqrt{\pi}[(\sqrt{d}+\log(2))^{2}+\log(2)]}$ and $2^{l^{*}-1}\alpha\leq\|z_{1}-z_{2}\|_{2}\leq 2^{l^{*}}\alpha$. Jensen’s inequality together with $\sqrt{1-t}\leq 1-t/2$ for $t\leq 1$ allow us to conclude.
Lemma E.2.
Let $\Omega$ be closed and convex and $S,S^{\prime}\in\cup_{m\geq 1}\Omega^{m}$ be two datasets, $V=\operatorname{Variation\_API}(S^{\prime})$ and $\hat{\mu}=\operatorname{NN\_histogram}(S,V,\rho)$ (see Algorithm 4). Suppose that for any $z_{1}\in\Omega,z_{2}\in S$, $\mathbb{E}[\min_{z\in\operatorname{Variation\_API}(z_{1})}\rho(z,z_{2})]\leq(1-\gamma)\rho(z_{1},z_{2})+\alpha$. Then,
@@
$\mathbb{E}[W_{1}(\mu_{S},\hat{\mu}_{t+1})]$ $=\mathbb{E}\left[\inf_{\pi\in\Pi(\mu_{S},\hat{\mu}_{t+1})}\int_{\Omega\times\Omega}\rho(z_{1},z_{2})d\pi(z_{1},z_{2})\right]$
@@
$=(1-\gamma)\int_{\Omega\times\Omega}\rho(z_{1},z_{2})d\pi^{\prime}(z_{1},z_{2})+\alpha,$
@@
$\mathbb{E}[W_{1}(\mu_{S},\hat{\mu}_{t+1})]\leq(1-\gamma)\int_{\Omega\times\Omega}\rho(z_{1},z_{2})d\pi^{\prime}(z_{1},z_{2})+\alpha.$
@@
E.2 Upper bounds
We need upper bounds on Gaussian complexity terms and on the convergence of empirical measures in $W_{1}$. Let us start showing an upper bound on the Gaussian complexity via Dudley chaining. This proof follows closely the proof of upper bounds for Laplacian Complexity given in [Vershynin23DPsyntheticdata].
Lemma E.4 (Gaussian Complexity bound).